OpenStudy (anonymous):

Can someone help factor this? a^3^y+1? I've been trying and I just can't find the right answer :(

OpenStudy (hoblos):

first is it \[a ^{3^{y}}\] or\[(a ^{3})^{y}\]

OpenStudy (anonymous):

the first one

OpenStudy (hoblos):

!! i dont think that it can be factorized but the second form can be factorized check it again

OpenStudy (anonymous):

It's definitely written down as the first one. but if oyu can help me iwth the second one I'll see if they'll take it

OpenStudy (hoblos):

ok (a^3)^y+1 = a^3y+1 = (a^y +1)(a^2y - a^y +1)

OpenStudy (anonymous):

It took that answer! Thank you so much!

OpenStudy (hoblos):

good XD glad to help =)